A 60-Ω electrical appliance and a 110-Ω electrical appliance are plugged into the outlets of a house. By what factor will the power used change if an additional 130-Ω electrical device is plugged in?
Answer #1

R1 = 60 II 110 = 38.82 ohm ( both are in parallel connections )

P1 = V^2/R1 = V^2/38.82 ( both are in parallel connections )

Now R2 = 60 II 110 II 130 = 29.895 ohm

P2 = V^2/R2 = V^2/29.895

Then P2/P1 = 38.82/29.895 = 1.298

feel free to drop comment if you have any doubt.